OpenStudy (anonymous):

\[\int\limits_{0}^{4} (3x+x^{3}/2) dx\]

OpenStudy (unklerhaukus):

\[3\int_0^4xdx+{1 \over 2}\int_0^4x^3dx\]

OpenStudy (unklerhaukus):

\[=3 { x^2 \over 2} \Big|_0^4+{1 \over 2}x^4/4\Big|_0^4 \]

OpenStudy (unklerhaukus):

\[=3\times 16/2+{1 \over2 } 256/4\] \[=24+32\] \[=56\]

OpenStudy (anonymous):

thanks a lot !